# Hunting for a 17 MeV particle coupled to electrons

### v2

@@ Abstract
We discuss a set of precision observables that can probe the existence of a light particle $X$ coupled to electrons in the mass range of 1–100 MeV. As a case study, we consider the recent excess of $e^{+}e^{-}$ final-state events at $\sqrt{s}=16.9$ MeV reported by the PADME collaboration. Interestingly, this mass is tantalizingly close to the invariant mass at which anomalous $e^{+}e^{-}$ pair production has previously been observed in nuclear transitions from excited to ground states by the ATOMKI collaboration. For the scenario in which the new particle has a vector coupling to electrons, we show that the PADME excess is already in tension with constraints from the anomalous magnetic moment of the electron and the non-observation of the exotic decay $\pi^{+}\to e^{+}\nu X$ at the SINDRUM experiment. Further improvements in the measurement of the electron $g$-2, together with upcoming results from PIONEER (searching for $\pi^{+}\to e^{+}\nu X$) and Mu3e (searching for $\mu^{+}\to e^{+}\overline{\nu}_{\mu}\nu_{e}X$), are expected to definitively probe this scenario in the near future. We also explore alternative possibilities where the new particle has scalar, pseudoscalar, or axial-vector couplings.
@@ I Introduction
In this work, we examine a set of observables sensitive to a new light particle coupled to electrons in the 1–100 MeV mass range. While exploring in more generality various coupling scenarios (scalar, pseudoscalar, vector, and axial-vector) we focus in particular on a 17 MeV state with vector couplings to electrons, motivated by the recent PADME analysis. Specifically, we assess current constraints and future sensitivities from the electron anomalous magnetic moment, exotic pion decays at the SINDRUM experiment Eichler et al. (1986), as well as upcoming results from PIONEER Altmannshofer et al. (2022) and Mu3e Arndt et al. (2021), which aim to search for exotic decays of the pion and muon, respectively. These experiments are expected to decisively test the above scenario in the near future.
@@ II Phenomenological setup
where $X=S,P,V,A$, and $\Gamma_{S}=1$ if $S$ is a scalar, $\Gamma_{P}=i\gamma_{5}$ if $P$ is a pseudoscalar, $\Gamma_{V}=\gamma_{\mu}$ if $V_{\mu}$ is a vector, and $\Gamma_{A}=\gamma_{\mu}\gamma_{5}$ if $A_{\mu}$ is an axial-vector.
@@ III Electron $\boldsymbol{g-2}$
[…]
@@
Therefore, $a_{e}$ can be used as a precision test of the SM and its extensions Giudice et al. (2012); Crivellin et al. (2018); Erdelyi et al. (2025). The contribution of the boson $X$ to the electron $g$-2 stemming from the Lagrangian of Eq. (1) reads Jegerlehner (2017); Athron et al. (2021):
@@
Comparing the results of Eqs. (11-14) with Eqs. (5) and (6) we learn that the PADME result, see Eq. (2), is in a significant tension with the electron $g$-2 bound. Hereafter, for definiteness, we assume three benchmark scenarios: 1) $|\Delta a_{e}|\leq 10^{-12}$, where we inflated the current experimental errors in $\delta a^{\alpha(Cs)}_{e}$ and $\delta a^{\alpha(Rb)}_{e}$ to make Eqs. (5) and (6) consistent, 2) $|\Delta a_{e}|\leq 10^{-13}$, assuming a resolution of the current discrepancy in the measurements of $\alpha$ with a precision of $\mathcal{O}(10^{-13})$, and 3) $|\Delta a_{e}|\leq 10^{-14}$, which is the ultimate expected uncertainty on $\Delta a_{e}$ if both $\delta a^{\rm exp}_{e}$ and $\delta a^{\alpha}_{e}$ will improve by roughly one order of magnitude Di Luzio et al. (2025).
@@ IV Pion decay
Another way to detect the $X$ boson is through the charged pion decay $\pi^{+}\to e^{+}\nu X$ [^3] For a related analysis, including also vector and axial-vector couplings to quarks, see Ref. Hostert and Pospelov (2023). The main focus here is on the electron coupling and the interplay of this observable with the PADME excess. The SINDRUM experiment searched for $e^{+}e^{-}$ resonances in $\pi^{+}\to e^{+}\nu X$, $X\to e^{+}e^{-}$ with sensitivity to branching ratios of $\mathcal{O}(10^{-10})$ Eichler et al. (1986). In order to constrain the parameter space of our scenarios, we exploit the bounds by SINDRUM Eichler et al. (1986) given as a function of the $X$ mass and lifetime. If $X$ is invisible, a complementary limit can be derived from a search for $\pi^{+}\to e^{+}\nu X$ at PIENU Aguilar-Arevalo et al. (2021). However, in the parameter space of our interest, we find that $X$ decays promptly into the visible $e^{+}e^{-}$ channel and therefore the PIENU bounds do not apply in our case. In the future, charged pion decays into visible final states can be searched for at the proposed experiment PIONEER Altmannshofer et al. (2022), planning to reach a sensitivity to branching ratios of $\mathcal{O}(10^{-11})$.
@@
When $X=S,P$, we find that $\Gamma(\pi^{+}\to\mu^{+}\nu X)\sim g^{2}_{eX}G^{2}_{F}m^{3}_{\pi}f^{2}_{\pi}|V_{ud}|^{2}/(4\pi)^{3}$, where $G_{F}$ is the Fermi constant, $f_{\pi}$ the pion decay constant, $V_{ud}$ the relevant CKM matrix element, and $1/(4\pi)^{3}$ stands for the three body phase space. Therefore, we find that
@@
where we used the standard pion decay rate $\Gamma(\pi^{+}\to\mu^{+}\nu)\approx G^{2}_{F}m_{\pi}m^{2}_{\mu}f^{2}_{\pi}|V_{ud}|^{2}/4\pi$. The current experimental bound $\mathcal{B}(\pi^{+}\!\to\!e^{+}\nu X)\lesssim 6\times 10^{-10}$ Eichler et al. (1986) by SINDRUM is therefore expected to translate into a bound on the coupling $g_{eX}\lesssim\,\text{few}\times 10^{-4}$.
On the other hand, if $X=V,A$ is a gauge boson coupled to a non-conserved current in an effective field theory where gauge invariance is restored at a scale $\Lambda$, its longitudinal mode can participate in processes at energies $E<\Lambda$, leading to enhancements by factors of $\mathcal{O}(E^{2}/m_{X}^{2})$. In the context of radiative pion decays, this implies an additional enhancement of $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$, which places even tighter constraints on light spin-1 mediators Dror et al. (2017); Dror (2020); Hostert and Pospelov (2023). [^4] We note that in the case of a conserved current, such as for a dark photon, the enhancement discussed above is absent, and the resulting bounds from pion decay are expected to be comparatively weaker. Although such a scenario is not captured by the simplified model in Eq. (1), the introduction of additional couplings of $X$ to other SM fermions would require the inclusion of further observables.
@@
when $X=S,P$, in good agreement with NDA in Eq. (15). Similarly, when $X=V,A$, we find
$\!\!\frac{\mathcal{B}(\pi^{+}\to e^{+}\nu X)}{6\times 10^{-10}}$ $\approx 87\left(\frac{g_{eX}}{5.6\!\times\!10^{-4}}\right)^{\!2}\,,$ (17)
again in good agreement with NDA after taking into account the enhancement factor $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ discussed above.
@@ V Muon decay
In comparison to pion decay experiments such as SINDRUM and PIONEER, the projected sensitivity of Mu3e is consistently weaker. This difference arises from the chirality suppression present in the SM background $\pi^{+}\to e^{+}\nu_{e}\gamma^{*},\gamma^{*}\to e^{+}e^{-}$, which does not apply to muon decays.
@@ VI Results
Figure 1: Present constraints (solid lines) and expected sensitivities (dashed lines) for a light $X=S,P,V,A$ particle coupled to electrons in the 1–100 MeV mass range. The vertical, red line at $m_{X}=16.9$ MeV denotes the $X_{17}$ benchmark, with the red star (in the vector case) representing the PADME best-fit value $g_{eV}=5.6\times 10^{-4}$.
@@
When the light particle is a spin-0 state, we find the limits to be comparable between the scalar ($X=S$ in the upper-left panel of Fig. 1) and the pseudoscalar ($X=P$ in the upper-right panel). In both cases, the most stringent constraints arise from measurements of the electron $g$-2 (dashed lines in different shades of green), with the precision in $|\Delta a_{e}|$ of $\mathcal{O}(10^{-14})$ allowing to probe the (pseudo)scalar couplings at the level of $g_{eX}\lesssim 10^{-5}$. A more moderate precision of $|\Delta a_{e}|\simeq\mathcal{O}(10^{-13})$ results in bounds which overlap to a good extent with the projections at the proposed PIONEER experiment which aims to reach sensitivities in radiative pion decays at the level of $\mathcal{B}(\pi^{+}\to e^{+}\nu X)\lesssim\mathcal{O}(10^{-11})$ (dashed orange line). Interestingly, the current bounds from the SINDRUM experiment (solid orange line) searching for the neutral particle emitted in $\pi^{+}$ decay approximately correspond to the sensitivity in $|\Delta a_{e}|\simeq\mathcal{O}(10^{-12})$ in the mass region above 40 MeV. At present, these constitute the most sensitive probes of new physics scenarios coupled to electrons in the 1–100 MeV mass range. Future muon experiments are expected to reach a sensitivity of $g_{eX}\lesssim 10^{-4}$ through the peak search in the invariant mass spectrum of $e^{+}e^{-}$ pair originating from a narrow resonance $X$, supplemented by the angular observables constructed from the $e^{\pm}$ momenta. The projected limits from the Mu3e experiment are shown as dashed blue lines, accounting for two planned phases: Phase I, with $\mathcal{O}(2.5\times 10^{15})$ muon decays, and Phase II, with an expected total of $\mathcal{O}(5.5\times 10^{16})$ events.
@@
When the light particle is a spin-1 state, the limits are overall more stringent than in the spin-0 case. This is particularly evident for the constraints from exotic $\pi^{+}$ decays, as set by SINDRUM (solid orange) and projected PIONEER (dashed orange) experiments. These stronger bounds arise due to the additional enhancement from the emission of the longitudinal mode, which contributes a factor of $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ to $\mathcal{B}(\pi^{+}\to e^{+}\nu X)$. This enhancement is especially significant in the low-mass region when compared with the spin-0 scenarios. Indeed, these pion decay limits represent the most stringent constraints, followed by those from projections on the precision of $\Delta a_{e}$ (dashed green lines), reversing the pattern observed in the spin-0 case. A similar trend holds for experiments utilizing intense muon beams: in Phase II of the Mu3e experiment, the spin-1 scenario is expected to be probed down to $g_{eX}\simeq 8\times 10^{-5}$ at the benchmark mass of $m_{X}=17$ MeV, compared to $g_{eX}\simeq 10^{-4}$ in the spin-0 case. Furthermore, the axial-vector scenario ($X=A$, shown in the lower-right panel of Fig. 1) is more tightly constrained by the electron $g$-2 than the vector case ($X=V$, shown in the lower-left panel). Remarkably, the PADME best-fit value of $g_{eV}=5.6\times 10^{-4}$ (indicated by the red star in Fig. 1) is already in tension with these constraints, and upcoming results from Mu3e and PIONEER will be well positioned to robustly test this scenario.

### v3

@@ Abstract
We discuss a set of precision observables that can probe the existence of a light particle $X$ coupled to electrons in the mass range of 1–100 MeV. As a case study, we consider the recent excess of $e^{+}e^{-}$ final-state events at $\sqrt{s}=16.9$ MeV reported by the PADME collaboration. Interestingly, this mass is tantalizingly close to the invariant mass at which anomalous $e^{+}e^{-}$ pair production has previously been observed in nuclear transitions from excited to ground states by the ATOMKI collaboration. For the scenario in which the new particle has a vector coupling to electrons, we show that the PADME excess is already in tension with constraints from the anomalous magnetic moment of the electron. Further improvements in the measurement of the electron gg-2, together with upcoming results from PIONEER (searching for $\pi^{+}\to e^{+}\nu X$) and Mu3e (searching for $\mu^{+}\to e^{+}\overline{\nu}_{\mu}\nu_{e}X$), are expected to definitively probe this scenario in the near future. We also explore alternative possibilities where the new particle has scalar, pseudoscalar, or axial-vector couplings.
@@ I Introduction
In this work, we examine a set of observables sensitive to a new light particle coupled to electrons in the 1–100 MeV mass range. While exploring in more generality various coupling scenarios (scalar, pseudoscalar, vector, and axial-vector) we focus in particular on a 17 MeV state with vector couplings to electrons, motivated by the recent PADME analysis. Specifically, we assess current constraints and future sensitivities from the electron anomalous magnetic moment, exotic pion decays at the SINDRUM experiment Eichler et al. (1986); Egli et al. (1989), as well as upcoming results from PIONEER Altmannshofer et al. (2022) and Mu3e Arndt et al. (2021), which aim to search for exotic decays of the pion and muon, respectively. These experiments are expected to decisively test the above scenario in the near future.
@@ II Phenomenological setup
where $X=S,P,V,A$, and $\Gamma_{S}=1$ if $S$ is a scalar, $\Gamma_{P}=i\gamma_{5}$ if PP is a pseudoscalar, $\Gamma_{V}=\gamma_{\mu}$ if $V_{\mu}$ is a vector, and $\Gamma_{A}=\gamma_{\mu}\gamma_{5}$ if $A_{\mu}$ is an axial-vector.
@@ III Electron $\boldsymbol{g-2}$
showing a disagreement of $5.5\,\sigma$. Using the above determinations of $\alpha$ to predict the Standard Model (SM) value $a^{\rm SM}_{e}$ and comparing it with the latest experimental measurement of $a^{\rm exp}_{e}=(115\,965\,218\,059~\pm~13)\times 10^{-14}$ Fan et al. (2023), yields the following values of $\Delta a_{e}\equiv a_{e}^{\rm exp}-a_{e}^{\rm SM}$:
@@
Therefore, $a_{e}$ can be used as a precision test of the SM and its extensions Giudice et al. (2012); Crivellin et al. (2018); Erdelyi et al. (2025). The contribution of the boson $X$ to the electron gg-2 stemming from the Lagrangian of Eq. (1) reads Jegerlehner (2017); Athron et al. (2021):
@@
Comparing the results of Eqs. (11-14) with Eqs. (5) and (6) we learn that the PADME result, see Eq. (2), is in a significant tension with the electron gg-2 bound. Hereafter, for definiteness, we assume three benchmark scenarios: 1) $|\Delta a_{e}|\leq 10^{-12}$, where we inflated the current experimental errors in $\delta a^{\alpha(Cs)}_{e}$ and $\delta a^{\alpha(Rb)}_{e}$ to make Eqs. (5) and (6) consistent, 2) $|\Delta a_{e}|\leq 10^{-13}$, assuming a resolution of the current discrepancy in the measurements of $\alpha$ with a precision of $\mathcal{O}(10^{-13})$, and 3) $|\Delta a_{e}|\leq 10^{-14}$, which is the ultimate expected uncertainty on $\Delta a_{e}$ if both $\delta a^{\rm exp}_{e}$ and $\delta a^{\alpha}_{e}$ will improve by roughly one order of magnitude Di Luzio et al. (2025).
@@ IV Pion decay
Another way to detect the $X$ boson is through the charged pion decay $\pi^{+}\to e^{+}\nu X$. [^3] For a related analysis, including also vector and axial-vector couplings to quarks, see Ref. Hostert and Pospelov (2023). The main focus here is on the electron coupling and the interplay of this observable with the PADME excess. The SINDRUM experiment searched for $e^{+}e^{-}$ resonances in $\pi^{+}\to e^{+}\nu X$, $X\to e^{+}e^{-}$ with sensitivity to branching ratios of $\mathcal{O}(10^{-10})$ Eichler et al. (1986); Egli et al. (1989). In order to constrain the parameter space of our scenarios, we exploit the bounds by SINDRUM Egli et al. (1989) given as a function of the $X$ mass. [^4] The signal in the search was modeled using the differential decay rate for a light scalar, whose kinematic features closely resemble those of the longitudinal modes of vector bosons. As shown in Hostert and Pospelov (2023), the small kinematic differences between the two cases are not expected to impact the interpretation of the resulting limits on spin-1 particles. If $X$ is invisible, a complementary limit can be derived from a search for $\pi^{+}\to e^{+}\nu X$ at PIENU Aguilar-Arevalo et al. (2021). However, in the parameter space of our interest, we find that $X$ decays promptly into the visible $e^{+}e^{-}$ channel and therefore the PIENU bounds do not apply in our case. In the future, charged pion decays into visible final states can be searched for at the proposed experiment PIONEER Altmannshofer et al. (2022), planning to reach a sensitivity to branching ratios of $\mathcal{O}(10^{-11})$.
@@
When $X=S,P$, we find that $\Gamma(\pi^{+}\to e^{+}\nu X)$ $\sim g^{2}_{eX}G^{2}_{F}m^{3}_{\pi}f^{2}_{\pi}|V_{ud}|^{2}/(4\pi)^{3}$, where $G_{F}$ is the Fermi constant, $f_{\pi}$ the pion decay constant, $V_{ud}$ the relevant CKM matrix element, and $1/(4\pi)^{3}$ stands for the three body phase space. Therefore, we find that
@@
where we used the standard pion decay rate $\Gamma(\pi^{+}\to\mu^{+}\nu)\approx G^{2}_{F}m_{\pi}m^{2}_{\mu}f^{2}_{\pi}|V_{ud}|^{2}/4\pi$. The current experimental bound $\mathcal{B}(\pi^{+}\!\to\!e^{+}\nu X)\lesssim 6\times 10^{-10}$ Eichler et al. (1986); Egli et al. (1989) by SINDRUM is therefore expected to translate into a bound on the coupling $g_{eX}\lesssim\,\text{few}\times 10^{-4}$.
@@
when $X=S,P$, in good agreement with NDA in Eq. (15).
On the other hand, if $X=V,A$ is a gauge boson coupled to a non-conserved current in an effective field theory where gauge invariance is restored at a scale $\Lambda$, its longitudinal mode can participate in processes at energies $E<\Lambda$, leading to enhancements by factors of $\mathcal{O}(E^{2}/m_{X}^{2})$.
This can be readily illustrated by considering the amplitude for $X$ emission in $\pi^{+}\to e^{+}\nu X$, arising from the effective Lagrangian
$\mathcal{L}$ $\supset\sqrt{2}G_{F}f_{\pi}V_{ud}^{*}\,(\partial_{\mu}\pi^{+})\,\overline{\nu}_{e}\gamma^{\mu}{\rm{P}_{L}}e$ (17)
$+g_{eX}X_{\mu}\left(Q_{e}\,\overline{e}\gamma^{\mu}{\rm{P}_{L}}e+Q_{\nu}\,\overline{\nu}_{e}\gamma^{\mu}{\rm{P}_{L}}\nu_{e}\right)\,$,
where $Q_{e}$ and $Q_{\nu}$ denote the $U(1)_{X}$ charges of the electron and electron neutrino, respectively. For the momentum configuration specified by $\pi^{+}(p_{1})\to e^{+}(p_{2})\nu_{e}(p_{3})X(p_{4})$, the amplitude reads
$\mathcal{M}$ $=\sqrt{2}G_{F}f_{\pi}V_{ud}^{*}\,g_{eX}\overline{u}(p_{3})\Gamma^{\mu}v(p_{2})\epsilon_{\mu}^{*}(p_{4})\,$, (18)
with
$\Gamma^{\mu}$ $=(Q_{e}-Q_{\nu})\gamma^{\mu}{\rm{P}_{L}}$ (19)
[…]
where $p_{ij}=p_{i}+p_{j}$.
In the case of weak-violation—where the components of an $SU(2)_{L}$ lepton doublet carry different charges under $U(1)_{X}$, the first term in Eq. (19) dominates the amplitude, leading to a large enhancement of $\mathcal{O}(m_{\pi}^{4}/m_{X}^{2}m_{e}^{2})$ in $\mathcal{B}(\pi^{+}\to e^{+}\nu X)$, and thus to strong constraints on light spin-1 mediators Dror et al. (2017); Dror (2020); Hostert and Pospelov (2023).
Importantly, this enhancement is absent in the weak-conserving case $Q_{e}=Q_{\nu}$, or more generally when the current is conserved, as in the case of a dark photon or $U(1)_{X=B-L}$ leading to significantly weaker bounds. Although such scenarios are not captured by the simplified model in Eq. (1), their limits can be understood in terms of the $Q_{e}=Q_{\nu}$ case in Eq. (17). Moreover, introducing couplings of $X$ to additional SM fermions would necessitate the inclusion of further observables.
Therefore, when $X=V,A$ and $Q_{\nu}=0$, we find
$\!\!\frac{\mathcal{B}(\pi^{+}\to e^{+}\nu X)}{6\times 10^{-10}}$ $\approx 87\left(\frac{Q_{e}\,g_{eX}}{5.6\!\times\!10^{-4}}\right)^{\!2}\,$, (20)
for $m_{X}=17$ MeV, again in good agreement with NDA after taking into account the enhancement factor $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ discussed above. Finally, when $Q_{e}=Q_{\nu}$, we have that
$\!\!\frac{\mathcal{B}(\pi^{+}\to e^{+}\nu X)}{6\times 10^{-10}}$ $\approx 6.5\times 10^{-3}\left(\frac{Q_{e}\,g_{eX}}{5.6\!\times\!10^{-4}}\right)^{\!2},$ (21)
demonstrating the chiral suppression evident in the amplitude in Eq. (19).
@@ V Muon decay
In comparison to pion decay experiments such as SINDRUM and PIONEER, the projected sensitivity of Mu3e is generally weaker if $X$ couples only to electrons. This difference arises from the chirality suppression present in the SM background $\pi^{+}\to e^{+}\nu_{e}\gamma^{*},\gamma^{*}\to e^{+}e^{-}$, which does not apply to muon decays. However, if $X=V,A$ couples in a weak-conserving manner with $Q_{e}=Q_{\nu}$ in Eq. (17), or more generally to a conserved current, the same chiral suppression also applies to the signal in pion decays, and Mu3e exhibits significantly better sensitivity in comparison.
@@ VI Results
Figure 1: Present constraints (solid lines) and expected sensitivities (dashed lines) for a light $X=S,P,V,A$ particle coupled to electrons in the 1–100 MeV mass range. The vertical, red line at $m_{X}=16.9$ MeV denotes the $X_{17}$ benchmark, with the red star (in the vector case) representing the PADME best-fit value $Q_{e}\,g_{eV}=5.6\times 10^{-4}$. The burgundy (orange) lines illustrate the constraints from the exotic pion decays under the hypothesis that $X=V,A$ couples to a conserved (non-conserved) current.
@@
When the light particle is a spin-0 state, we find the limits to be comparable between the scalar ($X=S$ in the upper-left panel of Fig. 1) and the pseudoscalar ($X=P$ in the upper-right panel). In both cases, the most stringent constraints arise from measurements of the electron gg-2 (dashed lines in different shades of green), with the precision in $|\Delta a_{e}|$ of $\mathcal{O}(10^{-14})$ allowing to probe the (pseudo)scalar couplings at the level of $g_{eX}\lesssim 10^{-5}$. A more moderate precision of $|\Delta a_{e}|\simeq\mathcal{O}(10^{-13})$ results in bounds which overlap to a good extent with the projections at the proposed PIONEER experiment which aims to reach sensitivities in radiative pion decays at the level of $\mathcal{B}(\pi^{+}\to e^{+}\nu X)\lesssim\mathcal{O}(10^{-11})$ (dashed orange line). Interestingly, the current bounds from the SINDRUM experiment (solid orange line) searching for the neutral particle emitted in $\pi^{+}$ decay approximately correspond to the sensitivity in $|\Delta a_{e}|\simeq\mathcal{O}(10^{-12})$ in the mass region above 40 MeV. At present, these constitute the most sensitive probes of new physics scenarios coupled to electrons in the 1–100 MeV mass range. Future muon experiments are expected to reach a sensitivity of $g_{eX}\lesssim 10^{-4}$ through the peak search in the invariant mass spectrum of $e^{+}e^{-}$ pair originating from a narrow resonance $X$, supplemented by the angular observables constructed from the $e^{\pm}$ momenta. The projected limits from the Mu3e experiment are shown as dashed blue lines, accounting for two planned phases: Phase I, with $\mathcal{O}(2.5\times 10^{15})$ muon decays, and Phase II, with an expected total of $\mathcal{O}(5.5\times 10^{16})$ events.
@@
When the light particle is a spin-1 state, the limits are overall more stringent than in the spin-0 case. This is particularly evident for the constraints from exotic $\pi^{+}$ decays, as set by SINDRUM (solid orange) and projected PIONEER (dashed orange) experiments when $Q_{e}\neq Q_{\nu}$ in Eq. (17), and we show the case with $Q_{\nu}=0$. These stronger bounds arise due to the absence of the chiral suppression in the weak-violating or non-conserved current case (see Sec. IV), and the additional enhancement from the emission of the longitudinal mode, which contributes a factor of $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ to $\mathcal{B}(\pi^{+}\to e^{+}\nu X)$. This enhancement is especially significant in the low-mass region when compared with the spin-0 scenarios. Indeed, these pion decay limits represent the most stringent constraints, followed by those from projections on the precision of $\Delta a_{e}$ (dashed green lines), reversing the pattern observed in the spin-0 case. However, when $Q_{e}=Q_{\nu}$, the aforementioned enhancements are absent, and the decay remains chirally suppressed, resulting in significantly weaker constraints from exotic pion decays probed by the SINDRUM and PIONEER experiments, shown by the solid and dashed burgundy lines in Fig. 1, respectively.
Regarding the experiments utilizing intense muon beams: in Phase II of the Mu3e experiment, the spin-1 scenario is expected to be probed down to $g_{eX}\simeq 8\times 10^{-5}$ (for $Q_{e}=1$ and $Q_{\nu}=0$) [^5] For the $Q_{e}=Q_{\nu}$ case, the difference in the $g_{eX}$ limit is $\mathcal{O}(1)$ and is therefore omitted from Fig. 1 to avoid further clutter. at the benchmark mass of $m_{X}=17$ MeV, compared to $g_{eX}\simeq 10^{-4}$ in the spin-0 case. Furthermore, the axial-vector scenario ($X=A$, shown in the lower-right panel of Fig. 1) is more tightly constrained by the electron gg-2 than the vector case ($X=V$, shown in the lower-left panel).
Remarkably, the PADME best-fit value of $Q_{e}\,g_{eV}=5.6\times 10^{-4}$ (indicated by the red star in Fig. 1) is already in tension with $\Delta a_{e}$ and excluded by SINDRUM when $Q_{e}\neq Q_{\nu}$. In contrast, when the new physics is weak-conserving with $Q_{e}=Q_{\nu}$, or $X_{\mu}$ couples to a conserved current, the PADME best-fit coupling is still allowed, but clearly within reach of upcoming experiments shown in Fig. 1.
Although we focused on the vector case motivated by the recent PADME analysis, we have also examined scalar, pseudoscalar, and axial-vector couplings, which may be relevant for other $X_{17}$ scenarios or in other contexts. As further data from PADME and related experiments becomes available, the framework explored here will remain a useful tool in scrutinizing the $X_{17}$ interpretation. Although we focused on the vector case motivated by the recent PADME analysis, we have also examined scalar, pseudoscalar, and axial-vector couplings, which may be relevant for other $X_{17}$ scenarios or in other contexts. As further data from PADME and related experiments becomes available, the framework explored here will remain a useful tool in scrutinizing the $X_{17}$ interpretation.